A populations consists of the following five values: 2, 3, 5, 7, 10.

List all possible sample of size 3.


1
Expert's answer
2022-05-06T03:06:13-0400

The number of possible samples which can be selected without replacement is

"\begin{pmatrix}
 N \\
 n
\end{pmatrix}=\cfrac{N! } {n! \cdot(N-n)! }=\\
=\cfrac{5! } {3! \cdot2! }=\cfrac{4\cdot5}{2}=10."


All the possible samples of sizes n=3 wich can be drawn without replacement from the population:

"\{ (2,3,5), (2,3,7),(2,3,10),(2,5,7),(2,5,10),\\
(2,7,10),(3,5,7),(3,5,10),(3,7,10),(5,7,10)\}."